OpenStudy (mathmagician):

one side of the right triangle is 48. Find all pairs of integers that could correspond another two sides (hypotenuse and the other side) one side of the right triangle is 48. Find all pairs of integers that could correspond another two sides (hypotenuse and the other side) @Mathematics

OpenStudy (anonymous):

14, 50

OpenStudy (anonymous):

have to think more :)

OpenStudy (cwrw238):

52, 20

OpenStudy (anonymous):

yeah, right

OpenStudy (cwrw238):

64, 80

OpenStudy (cwrw238):

all numbers in ratio3 ;4:5, also 13:12:5 will give you a right angled triangle

OpenStudy (anonymous):

also 7:24:25

OpenStudy (mathmagician):

also 577 and 575

OpenStudy (anonymous):

also 195, 189

OpenStudy (anonymous):

the basic thing is: express hypotenuse as h and the side as a then, we get h^2-a^2=(h+a)(h-a)=48^2 now factorise 48^2 in all ways possible to get h and a as integer u will get all the sets

OpenStudy (mathmagician):

yeah already began :) quite a fun ask